OpenStudy (mathisfun13):

Find the distance from point A (4, 1 and 1 half) to point B (2 and 2 thirds, -3)

ganeshie8 (ganeshie8):

\(\large A (4, \ \ 1\frac{1}{2})\) \(\large B (2\frac{2}{3} , -3)\)

ganeshie8 (ganeshie8):

like this ?

OpenStudy (mathisfun13):

yes

ganeshie8 (ganeshie8):

first convert the improper fractions.

ganeshie8 (ganeshie8):

you knw how to convert a 'mixed fraction' to an 'improper fractions' ?

OpenStudy (mathisfun13):

3/2? and 5/3

ganeshie8 (ganeshie8):

\(1 \frac{1}{2} = 1 + \frac{1}{2} = \frac{3}{2}\)

ganeshie8 (ganeshie8):

\(2\frac{2}{3} = 2 + \frac{2}{3} = \frac{\color{red}{8}}{3}\)

OpenStudy (mathisfun13):

right you multiply not add them ok

ganeshie8 (ganeshie8):

ha

ganeshie8 (ganeshie8):

use the distance formula now

OpenStudy (mathisfun13):

so subtract -3 and 3/2?

ganeshie8 (ganeshie8):

yes !

ganeshie8 (ganeshie8):

\(A (4, \ \frac{3}{2})\) \(B (\frac{8}{3}, -3)\)

ganeshie8 (ganeshie8):

how 2.5 ?

ganeshie8 (ganeshie8):

still stuck in this ?